### v2

# DRAN: A Distribution and Relation Adaptive Network for Spatio-temporal Forecasting

Abstract
Accurate predictions of spatio-temporal systems are crucial for tasks such as system management, control, and crisis prevention. However, the inherent time variance of many spatio-temporal systems poses challenges to achieving accurate predictions whenever stationarity is not granted. In order to address non-stationarity, we propose a Distribution and Relation Adaptive Network (DRAN) capable of dynamically adapting to relation and distribution changes over time. While temporal normalization and de-normalization are frequently used techniques to adapt to distribution shifts, this operation is not suitable for the spatio-temporal context as temporal normalization scales the time series of nodes and possibly disrupts the spatial relations among nodes. In order to address this problem, a Spatial Factor Learner (SFL) module is developed that enables the normalization and de-normalization process in spatio-temporal systems. To adapt to dynamic changes in spatial relationships among sensors, we propose a Dynamic-Static Fusion Learner (DSFL) module that effectively integrates features learned from both dynamic and static relations through an adaptive fusion ratio mechanism. Furthermore, we introduce a Stochastic Learner to capture the noisy components of spatio-temporal representations. Our approach outperforms state-of-the-art methods on weather prediction and traffic flow forecasting tasks. Experimental results show that our SFL efficiently preserves spatial relationships across various temporal normalization operations. Visualizations of the learned dynamic and static relations demonstrate that DSFL can capture both local and distant relationships between nodes.
Index Terms:
Spatio-temporal forecasting, graph neural network, distribution adaptation, adaptive network.
I Introduction
Spatio-temporal systems, characterized by intricate spatial interactions among senors (nodes) and rich temporal dynamics, are prevalent in various fields e. g., physics [1], meteorology [2, 3, 4], power grids [5, 6, 7] and transportation [8, 9]. Such systems often entail a large number of nodes, with interactions among nodes varying over time [10, 11]. The complexity of these systems and the time variance makes it challenging to efficiently manage future system developments and decision-making, hence requiring more accurate spatio-temporal prediction [12].
In spatio-temporal forecasting, historical time series data associated with nodes are used to predict future observations [13, 14]. Although numerous effective methods, particularly those based on deep learning, have been developed [15, 16, 17], predictions remains challenging due to the time variance of the environment-system interaction. This time variance can manifest in several ways, e. g., distribution shifts [18, 19], changes in spatial relationships among nodes [20], and changes in the nature of noise [21].
Normalization and de-normalization mechanisms are commonly employed to address distribution shifts in time series forecasting, particularly in scenarios that show temporal variation without strong spatial interactions [22, 23, 24, 25]. Normalization is applied to the input data to align its distribution with that of the training set, while de-normalization maps the processed results back to the original temporal distribution. However, above methods, which apply normalization and de-normalization on instance [22] or temporal dimensions [25], may not be well-suited for processing spatio-temporal data using graph neural networks that rely on relational inductive bias. In fact, temporal normalization scales the time series of nodes using different scaling factors as the temporal content evolves with time variance, hence causing erroneous scaled spatial messages propagating between nodes. Moreover, instance-level normalization uses the mean and standard deviation of all nodes over the input time span, addressing only coarse spatial and temporal distribution shifts, which is inadequate for adapting to the complex distributions in spatio-temporal contexts. Therefore, it is necessary to develop approaches that achieve temporal normalization while preserving spatial patterns in spatio-temporal forecasting tasks.
Various methods have been proposed to capture time-varying spatial relationships, including learning inherent static spatial structures from data [26, 27, 28], generating dynamic dependencies based on input windows [29, 30], and jointly modeling static and dynamic components through weighted fusion [31] or loss-constrained fusion ratios [32]. However, these approaches often overlook the dynamic interplay between static and dynamic spatial components, relying on fixed fusion schemes. This can lead to suboptimal spatio-temporal representations, especially in scenarios where the relative importance of static and dynamic features varies over time. Therefore, there is a pressing need for methods capable of adapting to both static and dynamic relations.
In addition, spatio-temporal systems are inherently affected by various sources of uncertainty, including data inaccuracies, sensor noise, and sudden stochastic events that might change with time. Given that these uncertainties impact prediction performance, various efforts have been made to quantify them. Some studies perform interval prediction to provide confidence regions for the forecasting horizon [33, 34]. Liang et al. [35] utilize latent variables to capture the uncertainties in air quality data, enhancing the accuracy of predictions.
In this paper, we propose a Distribution and Relation Adaptive Network (DRAN) that dynamically adapts to temporal variations and assesses uncertain components influenced by noise to achieve more comprehensive representation learning in spatio-temporal systems. More specifically, we develop a Spatial Factor Learner (SFL) that incorporates temporal normalization and de-normalization stages to mitigate the impact of distribution shifts while preserving spatial dependencies. To better accommodate spatio-temporal relations, we introduce a Dynamic-Static Fusion Learner (DSFL) module that merges features learned from sample-specific dynamic and static relations using an adaptive fusion ratio mechanism. Additionally, we propose a Stochastic Learner based on a Variational Autoencoder (VAE) to estimate uncertainty. Experimental validations on weather and traffic systems demonstrate that our model outperforms state-of-the-art methods. The novelty of our method can be summarized as follows:
  • We introduce an SFL module that enables a normalization and de-normalization mechanisms for distribution adaptation in a spatio-temporal context. The SFL module reverses temporal normalization while ensuring that spatial distributions closely match those of the forecasting horizons.
  • Different from previous spatio-temporal relations adaptive methods which fuse dynamic and static features with a fixed ratio [36, 37, 31, 32], we propose a DSFL module to adaptively combine features from both dynamic and static relations with a gating mechanism, accommodating the changing relations.
  • We propose a framework to learn both deterministic and stochastic spatio-temporal representations, where a VAE-based Stochastic Learner is introduced to capture the noisy components of the representations.
The paper is organized as follows. Section II introduces the state of the art. Section III details our network architecture and the overall workflow. Section IV describes the experimental setups, including dataset descriptions, network configurations, training procedures, and baseline comparisons. Section V presents the numerical results for traffic flow and weather predictions, accompanied by an ablation study and visualizations of module impacts. The discussion is provided in Section VI, and the conclusion in Section VII.
Figure 1: The architecture of DRAN. (a) provides an overview of DRAN, which learns both the deterministic (grey) and stochastic (yellow) components of the spatio-temporal representations. In learning the deterministic components, a normalization and de-normalization process (green) is conducted using the Spatial Factor Learner (SFL) module to achieve distribution adaptation. The Dynamic-Static Fusion Learner (DSFL) module (orange) learns spatial dependencies from both dynamic and static perspectives and fuses them using an adaptive ratio. (b), (c), and (d) offer detailed views of the structures of the DSFL, SFL, and Stochastic Learner, respectively.
II Related Works
II-A Spatio-temporal forecasting
Spatio-temporal forecasting is a longstanding and important research task. Early approaches primarily employed statistical models such as Historical Average (HA) and ARIMA [38, 39, 40], which offer interpretability but struggle to capture complex spatio-temporal dependencies. With the advent of deep learning, various modules are developed to model temporal and spatial dependencies more effectively. For temporal modeling, recurrent and convolutional neural networks [41, 42, 43] show effectiveness in short-term forecasting, while transformer-based architectures and state space models [44] demonstrate strong performance in both short- and long-term forecasting. Recurrent networks update historical information through hidden states, whereas transformer-based methods use attention mechanisms to capture dependencies from time-point [45], patch [46], and frequency-domain perspectives [47, 48]. To model spatial dependencies, both explicit and implicit methods have been proposed. Explicit methods typically utilize graph-based structures to represent node-link relationships and facilitate message propagation [28]. In contrast, implicit methods encode spatial dependencies within network parameters or attention mechanisms [49, 50, 51]. Despite promising performance, many existing methods still struggle with capturing evolving spatial structures and often lack interpretability. To address these issues, decomposition-based approaches emerge as a promising direction. These methods decompose temporal signals into trend and seasonal components [52, 53], extract multi-frequency patterns [54], and learn multi-scale spatial relationships [55, 56]. In this work, we propose a framework that decomposes spatial relations into dynamic and static components while learning both deterministic and stochastic features.
II-B Static and dynamic Relation Learning
Spatio-temporal forecasting methods are designed to capture both the static and dynamic relations existing among nodes. Static relations are typically learned from training datasets using trainable embeddings and adjacency matrices to depict inter-node relationships [57, 16, 58]. Conversely, other studies [59, 60] construct dynamic networks by generating sample-specific relations, which can be derived either directly from each input time series based on node similarity [56, 61] or through the use of meta-learners that generate dynamic parameters [62, 63]. Additionally, some methods refine the original relationships with input features [20, 64]. Furthermore, certain approaches simultaneously learn static and dynamic relations and integrate the gathered information. For instance, Fang et al. [32] decompose spatial relations into static and dynamic components, using a min-max learning paradigm with hyperparameters to simultaneously learn both aspects. Similarly, Li et al. [31] employ two parallel Graph Convolutional Networks (GCNs) to learn static and dynamic information, which is then fused using a weighted sum. These methods typically fuse the static and dynamic features at a fixed ratio. In this work, we focus on the adaptive fusion of dynamic and static spatial relations.
II-C Adaptation to Distribution Shifts
Normalization methods utilize statistic properties to normalize and de-normalize time series, adapting to distribution shifts in systems. Commonly, these methods utilize the mean and variance of input time series for data normalization. RevIN [22] employs a learnable affine transformation to align the means and standard deviations of inputs with those of outputs, facilitating distribution removal and reconstruction. Non-stationary Transformer [23] argues that existing stationary methods remove excessive information from time series, which hampers the model’s ability to learn temporal dependencies. To address this, it introduces De-stationary Attention modules that aim to balance this trade-off. Deep Adaptive Input Normalization (DAIN) [19] implements linear and gated layers to learn adaptive scaling and shifting factors for normalization. ST-Norm [65] proposes temporal and spatial normalization modules, which separately refines the high-frequency component and the local component of features. Additionally, U-Mixer [66] employs covariance analysis to correct stationarity for each feature. EAST-Net [67] generates sequence-specific network parameters to adapt dynamically to events.
III Methodology
III-A Problem Definition
The observed variables of node $i$ at time step $t$ can be referred to as a $C$-dimensional vector $\bm{X}_{t,i}\in\mathbb{R}^{C}$. Each observation at time step $t$ is the result of a stochastic process, drawn from a conditioned distribution $\bm{X}_{t,i}\sim p_{t,i}(\bm{X}_{t,i}|\bm{X}_{t-1},\bm{X}_{t-2},\cdots,\bm{X}_{t-k},\cdots)$, where $\bm{X}_{t-k}\in\mathbb{R}^{N\times C}$ denotes the observations from $N$ nodes at time step $t-k$ in the spatio-temporal framework. The probability distribution $p_{t,i}$ varies over time and differs across nodes, indicating that distribution shifts occur when observations across different time periods exhibit distinct distributions, i.e.,
$D(p_{t_{k},i},p_{t_{l},i})>\delta,$
where $\delta>0$ is a small threshold, $D(\cdot,\cdot)$ is a distance function that estimates the discrepancy between distributions, and $t_{k}$ and $t_{l}$ refer to two different time steps. In this work, we utilize Gaussian kernel density estimation [68] for distribution estimation and use KL divergence as a distance function.
Algorithm 1 DRAN Framework
0:  Spatio-temporal dataset $Ds$, hyperparameter $\alpha$, lookback length $L$, horizon length $H$, max epoch
0:  model parameters $\theta$ of DRAN
1:  Initialize model parameters $\theta$ and node embedding $E_{\rm a}$
2:  for epoch = 1 to max epoch do
3:     for each batch $\textbf{{X}}_{t-L:t},\textbf{{X}}_{t:t+H}\in Ds$ do
4:        Compute $\mu_{\bm{X}}$, $\sigma_{\bm{X}}$
5:        Obtain $\bm{\textit{X}}_{\rm Norm}$ via Eq (1)
6:        $\bm{\textit{X}}_{\rm tem}\xleftarrow{}\rm{DestationaryAtt}(\bm{\textit{X}}_{\rm Norm},\mu_{\bm{X}},\sigma_{\bm{X}})$ via Eq (2-4)
7:        $\mu_{\rm spa}$, $\sigma_{\rm spa}\xleftarrow{}\rm{SFL}(\bm{\textit{X}},\bm{\textit{X}}_{\rm tem},\mu_{\bm{X}},\sigma_{\bm{X}})$
8:        Rescale according to Eq (5) and obtain $\bm{\textit{X}}_{\rm spa}$
9:        $\rm{\textit{X}}_{\rm D}\xleftarrow{}\rm{DSFL}(\bm{\textit{X}}_{\rm spa},\bm{\textit{E}}_{\rm a})$ via Eq (6-11)
10:        $\rm{\textit{X}}_{\rm S},\rm{\textit{X}}_{\rm rec}\xleftarrow{}\rm{StochasticLearner}(\bm{\textit{X}}_{\rm D})$ via Eq (12-15)
11:        Compute $\mathcal{L}$ via Eq (17) and optimize $\theta$
12:     end for
13:  end for
14:  return  model parameters $\theta$
The goal of the spatio-temporal forecasting task is to develop a model $\mathrm{F}$ that uses the historical observations of length $L$ of nodes $\bm{X}_{t-L:t-1}\in\mathbb{R}^{L\times N\times C}$ to predict the future $H$-step observations of nodes $\bm{X}_{t:t+H}\in\mathbb{R}^{H\times N\times C}$, where $\bm{X}_{t-L:t}\in\mathbb{R}^{L\times N\times C}$ refers to time series of observed variables of all the nodes ($\bm{X}_{t-L},\bm{X}_{t-L+1},\cdots,\bm{X}_{t-1}$). Due to the time variance in spatio-temporal data generation, learning a model $\mathrm{F}$ that effectively handles shifting distributions is challenging. The overall workflow of DRAN and the structure of its modules are depicted in Fig. 1 and detailed in Algorithm 1.
III-B Distribution Adaptation
The distribution of historical time series often diverges from that of future time series due to time variance. Fig. 2 (a) demonstrates the effectiveness of temporal normalization, which aligns the historical distribution more closely with the future distribution. This alignment facilitates easier learning of the projection from the input to the forecasting horizon. Consequently, temporal normalization is both an effective and necessary operation for spatio-temporal forecasting.
Figure 2: Temporal and spatial distribution changes. We use the probability density function for distribution estimation and visualization. The probability density function is fitted using Gaussian Kernel Density, where the probability density at point $\xi$ is given by $\hat{f}(\xi)=1/mh\sum_{i=1}^{m}1/{2\pi}{e}^{-(\xi-\xi_{i})^{2}/{2h^{2}}}$ for random variables $\{\xi_{1},\cdots,\xi_{m}\}$, where $h$ represents the parameter that controls the smoothness of the distribution and is set to 0.1 in this work. Panel (a) illustrates the temporal distribution shifts of the NYCBike2 dataset and demonstrates the effectiveness of temporal normalization. Panel (b) displays the spatial distribution changes in the forecasting networks. "Lookback" and "Horizon" refer to the distribution of historical and future time series, respectively. Arrows of different colors represent the transformation to the horizon distributions.
While previous works [23, 25] have addressed distribution shifts in multivariate time series forecasting, these methods are not suitable for time series with spatial relations. Direct application of these stationary strategies to spatio-temporal forecasting results in degraded performance. As shown in Fig. 2 (b), the direct application of normalization (colored in green) leads to spatial distributions deviating significantly from the spatial distribution of horizon windows. This is because they perform normalization within the time series of each node, leading to inconsistent scaling among nodes that are spatially interconnected. Therefore, we develop a framework that learns the temporal distribution shift on each node and preserves the spatial relations of nodes, thus maintaining the efficiency of spatial layers.
To learn the deterministic part of the inputs, the input time series is firstly filtered by removing high-frequency noise. We then normalize the input time series of each node using the temporal mean $\mu_{\bm{X}}\in\mathbb{R}^{1\times N\times C}$ and standard deviation $\sigma_{\bm{X}}\in\mathbb{R}^{1\times N\times C}$ calculated for each input window.
$\bm{X}_{\rm Norm}=(\bm{X}-\mu_{\bm{X}})/\sigma_{\bm{X}},$ (1)
where $\bm{X}_{\rm Norm}$ represents the normalized time series. To preserve essential temporal information, we incorporate temporal dependency learning through the de-stationary attention mechanism from the Non-stationary Transformer [23]. This mechanism utilizes de-stationary factors $\mu_{\rm tem}$ and $\sigma_{\rm tem}$, which are learned from a multilayer perceptron (MLP).
$\log\sigma_{\rm tem}$ $=$ ${\rm MLP}(\sigma_{\bm{X}},\ \bm{X}_{\rm Norm}),$ (2)
$\mu_{\rm tem}$ $=$ ${\rm MLP}(\mu_{\bm{X}},\bm{X}_{\rm Norm}),$ (3)
$\bm{X}_{\rm tem}$ $=$ ${\rm Softmax}(\sigma_{\rm tem}\bm{Q}\bm{K}^{\rm T}+\mu_{\rm tem})\bm{V}.$ (4)
In this process, $\bm{Q}$, $\bm{K}$, and $\bm{V}$ represent the query, key, and value of attention, respectively, each derived from the projection of the input $\bm{X}_{\rm Norm}$. The ${\rm Softmax}$ activation function is applied thereafter. To reverse the spatial patterns before inputting features into the spatial layers, we aim to de-normalize the features of nodes $\bm{X}_{\rm tem}$ to preserve the spatial relations. Since the temporal representation $\bm{X}_{\rm tem}$ is transformed by the De-stationary attention module, the input statistics $\mu_{\bm{X}}$ and $\sigma_{\bm{X}}$ cannot be directly applied. As illustrated in Fig. 1 (c), we employ the SFL module to generate de-normalization factors $\mu_{\rm spa}$ and $\sigma_{\rm spa}$. In detail, we learn the node-wise features of the original input time series $\bm{X}$ and the temporal representation $\bm{X}_{\rm tem}$ using a 1-dimensional convolution on the temporal dimension and a Linear layer. Additionally, the statistics $\mu_{\bm{X}}$ and $\sigma_{\bm{X}}$ are processed through a linear layer for feature dimension alignment. Subsequently, we utilize a MLP to integrate node-wise features of the input, its temporal representation, and input data statistics to generate the de-normalization factors. Finally, spatial factors $\mu_{\rm spa}$ and $\sigma_{\rm spa}$ are applied to de-normalize $\bm{X}_{\rm tem}$.
$\bm{X}_{\rm spa}=\frac{1}{\sigma_{\rm spa}}\bm{X}_{\rm tem}+\mu_{\rm spa},$ (5)
where $\bm{X}_{\rm spa}$ represents the de-normalized result of $\bm{X}_{\rm tem}$.
As illustrated in Fig. 2, our framework generates representations whose temporal and spatial distributions closely align with those of the forecasting horizon.
III-C Dynamic-Static Fusion Learner
In this section, we propose the DSFL module, which simultaneously captures dynamic and static spatial information and integrates them through an adaptive fusion mechanism. To capture static information for a given task, we employ a trainable task-adaptive node embedding, akin to the approach used in the Adaptive Graph Convolutional Recurrent Network (AGCRN) [57]. Accounting for temporal variations within historical windows, we employ an adaptive node embedding $\bm{E}_{\rm a}\in\mathbbm{R}^{L\times N\times C}$ to encode the static features of historical time series. Dynamic features are derived from the spatial similarity of each input data through spatial attention mechanisms.
$\bm{A}_{\rm Dy}$ $=$ $\bm{Q}_{\rm spa}\bm{K}_{\rm spa}^{\rm T},$ (6)
$\bm{X}_{\rm Dy}$ $=$ ${\rm Softmax}(\bm{A}_{\rm Dy})\bm{V}_{\rm spa},$ (7)
where $\bm{Q}_{\rm spa}$, $\bm{K}_{\rm spa}$, and $\bm{V}_{\rm spa}\in\mathbbm{R}^{L\times N\times C}$ represent the value, key, and query of spatial attention, respectively. $\bm{A}_{\rm Dy}$ represents the adjacency matrix capturing the learned dynamic spatial relations, and $\bm{X}_{\rm Dy}$ denotes the resulting dynamic features. Then, we extract static features through graph aggregation. In this process, we utilize a fully dense adjacency matrix, $\bm{A}_{\rm St}\in\mathbbm{R}^{L\times N\times N}$, derived from the adaptive node embedding $\bm{E}_{\rm a}$, to depict the spatial relations.
$\bm{A}_{\rm St}$ $=$ $\bm{E}_{\rm a}\bm{E}_{\rm a}^{\rm T},$ (8)
$\bm{X}_{\rm St}$ $=$ $\bm{A}_{\rm St}\bm{X}_{\rm spa}\bm{W},$ (9)
where $\bm{W}\in\mathbbm{R}^{C\times C}$ is the parameter matrix, and $\bm{X}_{\rm St}$ represents the features learned from static relations.
As the relationships between static and dynamic features change over time, we utilize a gating mechanism to integrate these features. The gate signal $\bm{z}$ is generated to control the balance between dynamic and static features.
$\bm{X}_{\rm Cat}$ $=$ ${\rm F_{\rm Cat}}(\bm{X}_{\rm Dy},\ \bm{X}_{\rm St}),$ (10)
$\bm{z}$ $=$ $\rm Sigmoid(\rm FC({\bm{X}_{\rm Cat}})),$ (11)
$\bm{X}_{\rm D}$ $=$ $\bm{z}\odot{\rm FC}(\bm{X}_{\rm Dy})+(\mathbf{1}-\bm{z})\odot{\rm FC}(\bm{X}_{\rm St}),$ (12)
where $\rm F_{\text{Cat}}$ conducts concatenation operation on the last dimension of dynamic features $\bm{X}_{\rm Dy}$ and static features $\bm{X}_{\rm St}$, generating the concatenated features $\bm{X}_{\rm Cat}\in\mathbbm{R}^{L\times N\times 2C}$. We generate the gate signal $\bm{z}$ by applying a FC layer to map $\bm{X}_{\rm Cat}$ from $\mathbbm{R}^{L\times N\times 2C}$ to $\mathbbm{R}^{L\times N\times C}$, followed by a Sigmoid activation function; here, FC denotes the fully connected layer. Finally, we employ Equation (12) to fuse the dynamic and static features, producing the final deterministic features $\bm{X}_{\rm D}$, which represent the final deterministic features learned from the input time series.
TABLE I: Datasets details
Attributes Weather NYCBike1 NYCBike2 NYCTaxi PeMS04 PeMS08
Start time 1/1/2012 1/4/2014 1/7/2016 1/1/2015 1/1/2018 1/7/2016
End time 31/12/2022 30/9/2014 29/8/2016 1/3/2015 28/2/2018 31/8/2016
Sample rate 1 hour 30 minutes 30 minutes 30 minutes 5 minutes 5 minutes
Training set 5,623 3,023 1,912 1,912 10,181 10,700
Testing set 1,607 864 546 546 3,394 3,566
Validation set 803 431 274 274 3,394 3,566
Node number 263 128 200 200 307 170
Feature number 1 2 2 2 1 1
Input length 24 19 35 35 12 12
Output length 12 1 1 1 12 12
III-D Stochastic Learner
After obtaining the deterministic features, we utilize both backward and forward VAEs to generate the stochastic parts of the historical and forecasting time series. Firstly, the deterministic features $\bm{X}_{\rm D}$ are fed into the latent layers $\text{F}_{\text{lat}}(\cdot)$ to obtain the mean $\mu_{\rm sto}$ and standard deviation $\sigma_{\rm sto}$ of $\bm{X}_{\rm D}$, thereby capturing the distribution of the input data. Then, we sample the latent features $\bm{z}_{\rm l}$ from the distribution $\mathcal{N}(\mu_{\rm sto},\sigma_{\rm sto}^{2})$. These latent features $\bm{z}_{\rm l}$ are then processed through the reconstruction layer $\text{F}_{\text{rec}}(\cdot)$ to map back to the stochastic components of the time series. The processes for the backward and forward Stochastic Learners are detailed below:
$\mu_{\mathrm{b,sto}},\sigma_{\mathrm{b,sto}}$ $=$ ${\mathrm{F}}_{\mathrm{b,lat}}(\bm{X}_{\mathrm{D}}),\ \bm{z}_{\mathrm{b,l}}\sim\mathcal{N}(\mu_{\mathrm{b,sto}},\sigma_{\mathrm{b,sto}}^{2}),$ (13)
$\mu_{\mathrm{f,sto}},\sigma_{\mathrm{f,sto}}$ $=$ ${\mathrm{F}}_{\mathrm{f,lat}}(\bm{X}_{\mathrm{D}}),\ \bm{z}_{\mathrm{f,l}}\sim{\mathcal{N}}(\mu_{\mathrm{f,sto}},\sigma_{\mathrm{f,sto}}^{2}),$ (14)
$\bm{X}_{\mathrm{rec}}$ $=$ ${\mathrm{F}}_{\mathrm{b,rec}}(\bm{z}_{\mathrm{f,l}}),$ (15)
$\bm{X}_{\mathrm{S}}$ $=$ ${\mathrm{F}}_{\mathrm{f,rec}}(\bm{z}_{\mathrm{f,l}}),$ (16)
where $\mu_{\mathrm{b,sto}}$ and $\sigma_{\mathrm{b,sto}}$ are the mean and standard deviation of the backward features, respectively, and $\mu_{\mathrm{f,sto}}$ and $\sigma_{\mathrm{f,sto}}$ are those of the forward features. ${\rm F}_{\mathrm{b,lat}}$ and ${\rm F}_{\mathrm{f,lat}}$ represent the backward and forward latent layers, respectively, while ${\rm F}_{\mathrm{b,rec}}$ and ${\rm F}_{\mathrm{f,rec}}$ represent the backward and forward reconstruction layers. $\bm{X}_{\mathrm{rec}}$ denotes the reconstruction of the historical time series, and $\bm{X}_{\mathrm{S}}$ refers to the stochastic parts of the forecasting time series. To ensure that the Stochastic Learner can effectively capture the input dynamic-related stochastic components, we impose constraints on the learned latent feature distribution and the reconstruction values.
The loss function $\mathcal{L}$ consists of three components: prediction error $\mathcal{L}_{\mathrm{pred}}$, reconstruction error $\mathcal{L}_{\mathrm{rec}}$, and the distribution error computed using KL divergence $\mathcal{L}_{\mathrm{kl}}$.
$\mathcal{L}$ $=$ $\mathcal{L}_{\rm pred}(\bm{X}_{t:t+H},\hat{\bm{X}}_{t:t+H})$ (17)
$+\alpha\mathcal{L}_{\rm rec}(\bm{X}_{t-L:t-1},\hat{\bm{X}}_{t-L:t-1})+\beta\mathcal{L}_{\rm kl},$
where $\alpha$ and $\beta$ are hyperparameters that balance the importance of various loss functions, and $\hat{\bm{X}}$ represents the time series generated by the neural network. Both the prediction error $\mathcal{L}_{\mathrm{pred}}$ and the reconstruction error $\mathcal{L}_{\mathrm{rec}}$ are computed using the Mean Absolute Error (MAE).
TABLE II: Baseline methods.
Task type Methods Task adaptive Dynamic adaptive
Time series forecasting Dual-stage Attention-based Recurrent Neural Network (DA-RNN) [69]
InfoTS [70]
AutoTCL [71]
Spatio-temoral forecasting Temporal Graph Convolutional Network (TGCN)[26]
Spatio-Temporal Graph Convolutional Network (STGCN)[27]
Graph Convolutional Gate Recurrent Unit (GCGRU) [28]
Adaptive Graph Convolutional Recurrent Network (AGCRN)[57]
Attention based Spatio-Temporal Graph Convolutional Networks (ASTGCN)[72]
Diffusion Convolutional Recurrent Neural Network (DCRNN)[73]
Spatio-Temporal Adaptive Embedding transformer (STAEformer)[30]
Spatio-Temporal Self-Supervised Learning (ST-SSL)[20]
Meta-Graph Convolutional Recurrent Network (MegaCRN) [74]
Regularized Graph Structure Learning (RGSL) [16]
Time-Enhanced Spatio-Temporal Attention Model (TESTAM) [75]
Memory-based Drift Adaptation network (MemDA) [76]
Ours
TABLE III: The selection of $\alpha$.
Weather NYCBike1 NYCBike2 NYCTaxi PeMS04 PeMS08
$\alpha$ MAE $\alpha$ MAE $\alpha$ MAE $\alpha$ MAE $\alpha$ MAE $\alpha$ MAE
0.30 0.765 1.50 5.115 1.00 4.903 1.50 10.961 0.50 19.283 0.50 13.884
0.50 0.719 3.50 5.073 1.50 4.875 2.50 11.039 0.70 18.614 0.70 13.879
0.80 0.705 5.50 5.158 2.50 4.914 3.50 10.864 0.90 18.585 0.90 13.933
1.00 0.660 7.50 5.057 3.50 4.827 4.50 11.074 1.00 18.561 1.00 13.726
3.00 0.766 9.50 5.159 4.50 4.847 5.50 11.030 2.00 18.615 1.20 13.696
$\beta$ MAE $\beta$ MAE $\beta$ MAE $\beta$ MAE $\beta$ MAE $\beta$ MAE
0.10 0.670 0.10 5.390 0.10 5.130 0.10 11.264 0.10 32.290 0.10 20.790
0.50 0.665 0.50 5.269 0.50 4.910 0.50 10.758 0.50 25.774 0.50 13.818
1.00 0.664 1.00 5.173 1.00 4.866 1.00 11.136 1.00 18.585 1.00 13.717
5.00 0.658 5.00 4.982 5.00 5.025 5.00 11.114 5.00 18.545 5.00 13.822
10.00 0.672 10.00 5.303 10.00 4.938 10.00 11.161 10.00 18.387 10.00 13.696
Selection ($\alpha,\beta$) (1.00, 5.00) ($\alpha,\beta$) (7.50, 5.00) ($\alpha,\beta$) (3.50, 1.00) ($\alpha,\beta$) (3.50, 0.50) ($\alpha,\beta$) (1.00,10.00) ($\alpha,\beta$) (1.00, 1.00)
After obtaining the deterministic component $\bm{X}_{\rm D}$ and the stochastic component $\bm{X}_{\mathrm{S}}$ of the time series, we input them into the decoder to map the features to the forecasting results. In this paper, the decoder comprises stacks of fully connected (FC) layers.
$\hat{\bm{X}}_{t-L:t}={\rm Decoder}({\rm Concatenate}[\bm{X}_{\rm D},\ \bm{X}_{\rm S}]).$ (18)
IV Experimental Setups
In this section, we detail our experimental setup, including the datasets used, the hyperparameter settings for our networks, the baseline comparisons, and the training process.
IV-A Datasets
We conduct spatio-temporal forecasting tasks on weather and traffic systems to predict temperature and traffic flows.
[…]
NYC datasets. We use traffic flow datasets for bikes and taxis in New York, which have been preprocessed by Ji et al. [20]. These datasets are segmented into three categories: NYCBike1, NYCBike2, and NYCTaxi, recording inflows and outflows of city bikes and taxis every 30 minutes. The NYCBike1 dataset spans from April 1st, 2014, to September 30th, 2014. The NYCBike2 dataset covers the period from July 1st, 2016, to August 29th, 2016, while the NYCTaxi dataset ranges from January 1st, 2015, to March 1st, 2015. The input and output setups are consistent with those described in Ji et al. [20]. In NYCBike1, we predict the inflow and outflow of 128 grids 30 minutes ahead using historical records of 9.5 hours, with a sequence length of 19. In NYCBike2 and NYCTaxi datasets, we utilize historical time series of 17.5 hours, with a sequence length of 35. The number of grids in NYCBike2 and NYCTaxi datasets is 200.
PeMS04 and PeMS08 datasets. The PeMS04 and PeMS08 datasets are subsets of the PeMS (PeMS Traffic Monitoring) dataset [78], which includes real-time traffic flow data collected from loop detectors on California highways. Specifically, PeMS04 contains traffic flow records from the San Francisco Bay Area, covering the period from January 1, 2018, to February 28, 2018. PeMS08 encompasses traffic data from July 1, 2016, to August 31, 2016. Both datasets are sampled at 5-minute intervals. In our forecasting task, we aim to predict traffic flow one hour ahead based on the past one-hour records. Both the input and output lengths for each prediction task are set to 12 time steps.
Application features of these datasets are shown in Table I.
IV-B Training Details
After the exploration stage, the averaged hyperparameters—assumed consistent across benchmarks—are presented as follows: The dimension $C$ of the adaptive node embedding is 80, consistent with STAEformer [30]. Two-layer MLPs with a hidden dimension of 64 are employed to generate $\mu_{\rm tem}$ and $\sigma_{\rm tem}$. In the SFL module, the Conv1d layers are configured with an input channel equal to the length of the lookback window $L$, an output channel of 1, a kernel size of 3, and "circular padding" as defined in the PyTorch package. The Linear layers map the feature dimension to a hidden dimension of 64. For the DSFL module, both the de-stationary attention and spatial attention layers are set to 3. Each de-stationary attention module follows the parameter setup of STAEformer [30], with 4 attention heads and a feed-forward dimension of 256. In the DSFL module, the feature dimensions of $\bm{X_{\rm Dy}}$ and $\bm{X}_{\rm St}$ are both set to 160, and the number of attention heads for $\bm{X}_{\rm Dy}$ is also set to 4. In the Stochastic Learner, the latent layers include 3 Linear layers with ReLU activation functions, mapping the features to 64 dimensions. The reconstruction part of the Stochastic Learner consists of 3 Linear layers followed by ReLU activation functions, which remap the feature dimension from 64 to 160. The decoder comprises 2 Linear layers that fuse the deterministic and stochastic representations to produce the target feature outputs. The balance hyperparameters $\alpha$ and $\beta$ are fine-tuned experimentally to account for the stochastic nature and uncertainties of the datasets. Table III presents the variation in prediction error corresponding to different values of $\alpha$, with the value minimizing the error selected as optimal. Moreover, the training process is conducted on the Adam optimizer with a learning rate of 0.001 and a batch size of 32. The number of training epochs is set to 100. We split the datasets into training, validation and test set with a ratio shown in Table I. Numerical experiments for the methods are conducted using various random seeds from the set $\{31,32,33,34,35\}$ to obtain the average performance and standard deviation.
IV-C Baselines
We compare our method against several baseline approaches, including state-of-the-art multi-variable time series forecasting methods and spatio-temporal forecasting methods. Spatio-temporal forecasting techniques can be classified into two categories based on how they learn spatial relations: task-adaptive and dynamic-adaptive methods. Task-adaptive methods focus on learning static temporal and spatial relations from training datasets, which remain fixed during the testing phase. In contrast, dynamic-adaptive methods capture dynamically changing relations from input windows or by updating memory with incoming data. Details of the baseline methods are provided in Table II. In multi-variable time series forecasting, DA-RNN is a classic dual-stage attention-based recurrent neural network designed to capture long-term temporal dependencies. InfoTS and AutoTCL focus on enhancing time series representation learning through series augmentations and contrastive learning. InfoTS introduces a novel contrastive learning approach with information-aware augmentations that adaptively select optimal augmentations and a meta-learner network to learn from datasets. AutoTCL achieves unified and meaningful time series augmentations at both the dataset and instance levels, leveraging information theory to enhance representation quality. TGCN, STGCN, DCRNN and GCGRU utilize physical spatial relations as adjacency matrices and employ static neural networks for prediction. TGCN, DCRNN, and GCGRU use Graph Convolutional Networks (GCN) and Recurrent Neural Networks (RNN) to capture spatial and temporal features. STGCN combines temporal and graph convolutions to learn spatial and temporal dependencies. AGCRN utilizes learnable node embeddings to adapt spatial relations to tasks and node-adaptive parameters to capture specific attributes of each node. ASTGCN and STAEformer employ attention mechanisms to capture dynamic changes in input features. RGSL learns spatio-temporal dependencies from a predefined graph and learnable node embeddings, dynamically fusing features from two graphs using an attention mechanism. Additionally, STAEformer employs learnable node embeddings and concatenates it with input time series to capture static features. MegaCRN utilizes node embeddings to learn the static relations and memory networks to dynamically match sample patterns with learned static features. ST-SSL does not learn static relations and only uses the adjacency matrix based on node distances as a prior graph. It fine-tunes the static graph using node similarities, effectively fusing dynamic and static features. TESTAM employs a mixture-of-experts model with three experts: one for temporal modeling, one for spatio-temporal modeling with a static graph, and one for spatio-temporal dependency modeling with a dynamic graph. We also fine-tune the hyperparameters of baseline methods to ensure fairness, especially for datasets not included in their official implementations. We evaluate the performance of these models using two metrics: MAE and Mean Absolute Percentage Error (MAPE). The number of samples in the test datasets is denoted as $m$. $\hat{\bm{X}}$ and $\bm{X}$ denote the predicted and actual observations of spatio-temporal systems.
${\mathrm{MAE}}=\frac{1}{m}\sum_{j=1}^{m}\left|\hat{\bm{X}}-\bm{X}\right|,$ (19)
${\mathrm{MAPE}}=\frac{100\%}{m}\sum_{j=1}^{m}\left|\frac{\hat{\bm{X}}-\bm{X}}{\bm{X}}\right|.$ (20)
V Experimental Results
V-A Comparison Results
As shown in Tables IV and V, numerical experiments, present the average performance and standard deviation of both the baseline methods and our DRAN model. Our DRAN model outperforms the baseline methods, demonstrating the best overall performance. DRAN achieves the best average performance on the Weather and NYC datasets, and ranks second on the PeMS04 and PeMS08 datasets. Our method performs less effectively on the MAPE metric but excels on the MAE metric. This suggests that the model prioritizes optimizing the overall error rather than minimizing relative error in specific scenarios, such as scenarios with small target values. Consequently, while DRAN’s performance on the MAPE metric is slightly lower than that of other methods, its superior MAE results demonstrate effective overall error control. In terms of prediction stability, DRAN exhibits slightly larger standard deviations compared to other methods across experiments with different random seeds. This is likely due to the model’s attempt to learn noise components in the representation, which are randomly sampled from distributions.
ST-SSL exhibits strong performance in one-step traffic flow prediction but is less effective on other datasets. This discrepancy may be due to ST-SSL’s emphasis on one-step ahead prediction, as its decoder is not optimized for multi-step forecasting. Additionally, MegaCRN performs well across both tasks, likely due to its task-adaptive node embeddings and the dynamic features learned from its meta memory networks. RGSL benefits from dynamically fusing an explicit prior graph with a learned implicit graph using attention mechanisms. STAEformer performs well on some datasets but poorly on others, which may be attributed to its non-adaptive fusion process. These results suggest that our proposed DRAN model is capable of adapting to various tasks and effectively capturing both static and dynamic features.
TABLE IV: The prediction results on weather and NYC datasets.
Methods & Metrics Datasets Weather NYCBike1 NYCBike2 NYCTaxi
Temperature Inflow Outflow Inflow Outflow Inflow Outflow
DA-RNN [69] MAE 7.951$\pm$1.036 14.158$\pm$3.523 14.565$\pm$3.383 13.192$\pm$1.014 13.115$\pm$1.178 24.642$\pm$7.398 28.721$\pm$8.953
MAPE (%) 2.709$\pm$0.355 60.359$\pm$12.515 62.955$\pm$12.885 53.894$\pm$4.686 60.253$\pm$13.219 67.319$\pm$9.888 70.737$\pm$14.281
InfoTS [70] MAE 1.343$\pm$0.031 6.436$\pm$0.133 6.797$\pm$0.138 6.641$\pm$0.091 6.087$\pm$0.096 14.912$\pm$0.239 11.545$\pm$0.225
MAPE (%) 0.459$\pm$0.011 33.724$\pm$0.553 34.920$\pm$0.551 32.231$\pm$0.634 30.113$\pm$0.603 21.770$\pm$0.442 20.982$\pm$0.409
AutoTCL [71] MAE 1.194$\pm$0.025 6.002$\pm$0.037 6.424$\pm$0.034 6.011$\pm$0.060 5.533$\pm$0.063 14.843$\pm$0.162 11.394$\pm$0.095
MAPE (%) 0.408$\pm$0.008 28.076$\pm$0.201 29.573$\pm$0.164 29.512$\pm$0.380 27.766$\pm$0.409 22.061$\pm$0.189 21.140$\pm$0.136
STGCN [27] MAE 2.278$\pm$1.103 17.106$\pm$0.016 17.177$\pm$0.216 32.981$\pm$35.614 30.180$\pm$28.265 78.624$\pm$40.578 65.907$\pm$33.232
MAPE (%) 0.778$\pm$0.377 58.208$\pm$1.926 58.789$\pm$1.383 63.402$\pm$19.134 62.237$\pm$13.177 84.124$\pm$30.571 75.094$\pm$24.967
TGCN [26] MAE 1.736$\pm$0.847 7.412$\pm$0.052 7.677$\pm$0.441 6.103$\pm$0.106 5.712$\pm$0.072 25.307$\pm$11.321 21.715$\pm$9.109
MAPE (%) 0.592$\pm$0.287 34.611$\pm$0.827 35.085$\pm$1.704 29.717$\pm$0.592 28.286$\pm$0.632 44.536$\pm$10.924 45.181$\pm$10.584
MemDA [76] MAE 1.786$\pm$0.071 6.777$\pm$0.154 7.246$\pm$0.100 6.452$\pm$0.127 6.284$\pm$0.252 18.505$\pm$0.661 17.199$\pm$0.542
MAPE (%) 0.611$\pm$0.026 29.344$\pm$0.543 30.351$\pm$0.126 29.312$\pm$0.571 27.995$\pm$0.802 28.719$\pm$1.861 32.463$\pm$1.726
ASTGCN [72] MAE 1.521$\pm$0.155 6.481$\pm$0.352 6.698$\pm$0.541 6.658$\pm$0.326 7.548$\pm$3.533 15.389$\pm$5.711 12.500$\pm$3.923
MAPE (%) 0.520$\pm$0.054 30.780$\pm$1.221 31.264$\pm$1.976 28.674$\pm$1.021 27.028$\pm$1.123 24.731$\pm$0.916 24.443$\pm$1.578
TESTAM [75] MAE 1.481$\pm$1.203 6.331$\pm$0.736 6.826$\pm$0.891 6.437$\pm$0.852 6.266$\pm$0.901 17.684$\pm$1.814 15.967$\pm$2.758
MAPE (%) 0.507$\pm$0.415 30.418$\pm$2.571 31.628$\pm$2.949 28.565$\pm$3.457 27.762$\pm$2.620 25.703$\pm$3.433 25.988$\pm$4.105
AGCRN [57] MAE 1.094$\pm$0.039 5.066$\pm$0.020 5.366$\pm$0.023 5.279$\pm$0.058 4.941$\pm$0.035 12.263$\pm$0.132 9.720$\pm$0.098
MAPE (%) 0.373$\pm$0.013 25.551$\pm$0.200 26.564$\pm$0.209 25.498$\pm$0.268 24.633$\pm$0.286 18.234$\pm$0.324 18.239$\pm$0.278
GCGRU [28] MAE 1.012$\pm$0.041 5.457$\pm$0.063 5.631$\pm$0.290 5.313$\pm$0.064 4.987$\pm$0.056 12.153$\pm$2.649 10.277$\pm$1.340
MAPE (%) 0.346$\pm$0.015 26.867$\pm$0.700 27.292$\pm$1.476 25.428$\pm$0.423 24.490$\pm$0.185 22.386$\pm$7.043 22.800$\pm$7.607
DCRNN [73] MAE 0.984$\pm$0.028 5.374$\pm$0.039 5.557$\pm$0.305 5.386$\pm$0.036 5.047$\pm$0.031 12.097$\pm$3.639 9.833$\pm$2.314
MAPE (%) 0.336$\pm$0.010 26.772$\pm$0.850 27.165$\pm$1.507 25.279$\pm$0.273 24.332$\pm$0.164 21.182$\pm$3.214 21.598$\pm$3.360
STAEformer [30] MAE 3.728$\pm$2.367 5.168$\pm$0.029 5.475$\pm$0.028 5.453$\pm$0.100 5.112$\pm$0.147 12.262$\pm$0.237 9.824$\pm$0.109
MAPE (%) 1.284$\pm$0.819 25.829$\pm$0.331 26.889$\pm$0.242 25.774$\pm$0.429 24.840$\pm$0.447 17.889$\pm$0.777 18.203$\pm$0.684
MegaCRN [74] MAE 0.952$\pm$0.027 5.042$\pm$0.016 5.357$\pm$0.043 5.152$\pm$0.049 4.826$\pm$0.021 12.084$\pm$0.158 9.569$\pm$0.175
MAPE (%) 0.325$\pm$0.010 25.329$\pm$0.167 26.275$\pm$0.213 24.878$\pm$0.409 23.870$\pm$0.293 17.883$\pm$0.204 17.664$\pm$0.338
RGSL [16] MAE 0.727$\pm$0.003 5.149$\pm$0.140 5.335$\pm$0.200 5.251$\pm$0.039 4.915$\pm$0.060 12.083$\pm$0.125 9.582$\pm$0.072
MAPE (%) 0.248$\pm$0.001 25.612$\pm$0.183 26.110$\pm$0.924 24.186$\pm$2.957 23.288$\pm$2.249 18.430$\pm$0.531 18.205$\pm$0.319
ST-SSL [20] MAE 1.352$\pm$0.010 5.135$\pm$0.024 5.265$\pm$0.023 5.042$\pm$0.029 4.714$\pm$0.023 12.010$\pm$0.481 9.790$\pm$0.101
MAPE (%) 0.461$\pm$0.004 25.430$\pm$0.295 24.605$\pm$0.265 22.633$\pm$0.112 21.813$\pm$0.808 16.383$\pm$0.100 16.855$\pm$0.228
Ours MAE 0.672$\pm$0.007 4.882$\pm$0.032 5.176$\pm$0.054 5.008$\pm$0.056 4.653$\pm$0.036 11.929$\pm$0.054 9.539$\pm$0.054
MAPE (%) 0.229$\pm$0.002 23.553$\pm$0.169 24.607$\pm$0.497 22.384$\pm$0.211 21.429$\pm$0.208 16.338$\pm$0.409 16.666$\pm$0.201
  • Results with bold are the overall best performance, and shading results have the suboptimal performance.
Figure 3: Prediction errors for the weather dataset. (a), (b), and (c) show the absolute prediction errors for DRAN, RGSL, and MegaCRN, respectively.
Figure 4: Prediction results for the NYCBike1 dataset. The city is partitioned into a grid map. Panels (a) and (f) in the first column show the actual bike flow, while panels (b) and (g) display the predictions made by our DRAN model. Panels (c) and (h) present the predictions from the sub-optimal method ST-SSL. Panels (d) and (i) illustrate the absolute prediction errors for DRAN, and panels (e) and (j) depict the absolute prediction errors for ST-SSL.
Figure 5: Visualization of temporal prediction results. (a) and (b) display the predicted temperature of Weather dataset of node 20 and node 190 from November 3rd, 2020 to November 17th, 2019. (c) and (d) show the predicted traffic inflow and outflow of NYCBike1 dataset of node 50 from 0:00 of August 25th, 2014 to 12:00 of September 1st.
Furthermore, we display the prediction results of our DRAN and the sub-optimal methods. As is shown in Fig. 3, by comparing the prediction errors of our DRAN, RGSL, and MegaCRN, we find that while the numerical metrics of these methods are very close, the distribution of prediction errors is different. The prediction results of our method demonstrate a more stable performance across all spatial regions, whereas other methods exhibit significantly larger errors in certain regions. This indicates that our method is more stable in prediction and better adapts to nodes with complex dynamic changes. As shown in Fig. 4, we provide two cases as examples to visualize the prediction results. We can see that both methods generate predictions that are similar to the ground truth. However, when comparing spatial prediction errors, DRAN shows fewer grids with large prediction errors. Additionally, we present the predicted time series of nodes in Fig. 5. In the weather dataset, both RGSL and our method capture the overall temporal trends of nodes and perform well in the more regular periodic variations, though they lack accuracy in some extreme values. This may be due to an insufficient ability to capture abrupt changes in the time series. In Fig. 5 (c) and (d), DRAN demonstrates superior performance in predicting the sudden decrease in traffic flow.
To balance computational cost and prediction accuracy, we compare inference times in Table VI, where methods are listed in descending order of inference time. The comparison is conducted using input data of the same size, repeated 100 times to compute the average inference time for a batch of NYCTaxi data. All experiments are performed on an Nvidia RTX3090 GPU. While DA-RNN achieves the fastest inference speed, it lacks sufficient prediction accuracy. DRAN delivers the best prediction accuracy with a moderate latency at inference time, comparable to ST-SSL and STAEformer.
V-B The Preservation of Spatial Distribution
To evaluate whether the SFL module preserves spatial distributions across various tasks or not, we analyze the spatial distributions of representations before SFL ($\bm{X}_{\rm tem}$), after SFL ($\bm{X}_{\rm spa}$), and at the forecasting horizon ($\bm{X}_{t:t+H}$). The objective is to determine whether the distributions of representations after SFL are closer to those at the forecasting horizon. For each forecasting task, we randomly generate indices and select samples from $\bm{X}_{\rm tem}$, $\bm{X}_{\rm spa}$, and $\bm{X}_{t:t+H}$. We then use Gaussian Kernel Density Estimation to assess the distributions of these representations. As visualized in Fig. 6, the results demonstrate the SFL module’s ability to preserve the spatial distribution of nodes. The primary goal of SFL is to align the spatial distribution of learned representations with that of the forecasting horizon, thereby enhancing accuracy in capturing spatial and temporal distribution changes. In Fig. 6, significant discrepancies between the spatial distribution of representations before SFL and the final horizon are observed. The SFL module effectively reduces these variations in the spatial distributions of learned representations, thereby facilitating the learning process to map spatially preserved representations to predictions.
TABLE V: The prediction results on PeMS04 and PeMS08 datasets.
Methods & Metrics Datasets PeMS04 PeMS08
Flow Flow
DA-RNN [69] MAE 104.741$\pm$12.184 92.276$\pm$9.643
MAPE (%) 137.640$\pm$18.494 66.774$\pm$13.187
InfoTS [70] MAE 25.851$\pm$0.510 24.006$\pm$1.225
MAPE (%) 19.556$\pm$0.464 14.682$\pm$0.666
AutoTCL [71] MAE 23.814$\pm$0.048 17.355$\pm$0.200
MAPE (%) 20.879$\pm$0.098 13.006$\pm$0.121
TGCN [26] MAE 34.785$\pm$0.204 24.448$\pm$0.129
MAPE (%) 27.972$\pm$0.754 14.589$\pm$0.219
GCGRU [28] MAE 24.398$\pm$0.043 17.778$\pm$3.356
MAPE (%) 15.105$\pm$0.464 10.622$\pm$2.838
STGCN [27] MAE 25.302$\pm$2.316 25.338$\pm$10.825
MAPE (%) 24.321$\pm$8.603 15.404$\pm$7.083
DCRNN [73] MAE 24.912$\pm$2.064 20.853$\pm$0.062
MAPE (%) 17.720$\pm$0.973 11.864$\pm$0.094
ASTGCN [72] MAE 23.565$\pm$0.942 20.195$\pm$0.398
MAPE (%) 16.813$\pm$1.004 11.303$\pm$0.552
ST-SSL [20] MAE 22.836$\pm$0.705 18.559$\pm$0.097
MAPE (%) 14.318$\pm$0.435 10.615$\pm$0.122
MemDA [76] MAE 20.037$\pm$0.150 16.370$\pm$0.207
MAPE (%) 11.969$\pm$0.100 9.342$\pm$0.243
RGSL [16] MAE 19.544$\pm$0.2571 17.452$\pm$3.102
MAPE (%) 13.862$\pm$0.242 9.186$\pm$0.182
TESTAM [75] MAE 19.331$\pm$0.481 15.757$\pm$0.360
MAPE (%) 12.098$\pm$0.419 9.083$\pm$0.215
AGCRN [57] MAE 19.3291$\pm$0.3053 16.101$\pm$0.085
MAPE (%) 12.937$\pm$0.035 9.219$\pm$0.065
MegaCRN [74] MAE 18.858$\pm$0.0413 15.597$\pm$0.245
MAPE (%) 12.808$\pm$0.098 8.834$\pm$0.096
STAEformer [30] MAE 18.241$\pm$0.082 13.538$\pm$0.039
MAPE (%) 12.064$\pm$0.071 8.858$\pm$0.017
Ours MAE 18.375$\pm$0.087 13.690$\pm$0.085
MAPE (%) 12.026$\pm$0.419 8.987$\pm$0.057
  • Results with bold are the overall best performance, and shading results have the suboptimal performance.
TABLE VI: Inference time comparison
Methods Time (s) MAE
GConvGRU [28] 1.383 12.022
DCRNN [73] 0.967 11.924
TGCN [26] 0.469 28.475
STGCN [27] 0.343 12.286
RGSL [16] 0.246 10.833
InfoTS [70] 0.179 13.228
MegaRCN [74] 0.178 10.827
AGCRN [57] 0.126 10.992
ASTGCN [72] 0.114 14.422
TESTAM [75] 0.075 16.826
DRAN (Ours) 0.075 10.734
ST-SSL [20] 0.073 10.996
STAEformer [30] 0.065 10.810
MemDA [76] 0.056 17.852
AutoTCL [71] 0.052 13.119
DA-RNN [69] 0.048 26.596
  • Experiments are conducted on a batch of NYCTaxi date which contains 8 samples. Results with bold are the fastest method and the method with best prediction performance.
TABLE VII: The effectiveness of SFL on various temporal normalization methods.
Weather NYCBike1 NYCBike2 NYCTaxi
   Strategies  Datasets Temperature Inflow Outflow Inflow Outflow Inflow Outflow
+RevIN MAE 0.732$\pm$0.004 5.031$\pm$0.064 5.341$\pm$0.094 5.192$\pm$0.068 4.831$\pm$0.060 12.301$\pm$0.217 9.714$\pm$0.168
+DAIN MAE 1.035$\pm$0.006 5.212$\pm$0.088 5.508$\pm$0.084 5.670$\pm$0.420 5.258$\pm$0.290 13.747$\pm$0.174 10.848$\pm$0.088
+DAIN+SFL MAE 0.663$\pm$0.003 4.942$\pm$0.050 5.229$\pm$0.028 5.268$\pm$0.135 4.899$\pm$0.159 12.303$\pm$0.147 9.828$\pm$0.171
+Non-st MAE 0.738$\pm$0.006 5.096$\pm$0.109 5.368$\pm$0.089 5.297$\pm$0.073 4.924$\pm$0.072 13.344$\pm$0.051 10.540$\pm$0.159
+Non-st+SFL MAE 0.671$\pm$0.006 4.882$\pm$0.032 5.176$\pm$0.054 5.008$\pm$0.056 4.653$\pm$0.036 11.929$\pm$0.054 9.539$\pm$0.054
+Dish-TS MAE 0.764$\pm$0.003 5.024$\pm$0.080 5.370$\pm$0.101 5.215$\pm$0.017 4.843$\pm$0.022 12.363$\pm$0.314 9.868$\pm$0.223
+Dish-TS+SFL MAE 0.676$\pm$0.008 4.998$\pm$0.094 5.317$\pm$0.079 5.077$\pm$0.050 4.777$\pm$0.053 12.208$\pm$0.208 9.720$\pm$0.133
+ST-norm MAE 1.288$\pm$0.038 5.205$\pm$0.069 5.469$\pm$0.072 9.782$\pm$0.526 9.455$\pm$0.500 13.993$\pm$0.621 11.252$\pm$0.529
+T-norm MAE 1.712$\pm$0.046 5.291$\pm$0.245 5.556$\pm$0.146 8.876$\pm$0.935 8.704$\pm$0.932 13.469$\pm$0.292 10.788$\pm$0.319
+T-norm+SFL MAE 0.747$\pm$0.165 4.952$\pm$0.056 5.239$\pm$0.031 8.479$\pm$0.211 8.204$\pm$0.240 12.149$\pm$0.124 9.654$\pm$0.055
Figure 6: The preservation of spatial distribution by the SFL module across the Weather, NYCBike1, NYCBike2, and NYCTaxi datasets is illustrated in panels (a), (b), (c), and (d). These panels showcase changes in the spatial distribution of latent representations and forecasting horizons. "Before SFL", "After SFL", and "Horizon" denote the spatial distribution of representations before the SFL module $\bm{X}_{\rm tem}$, after the SFL module $\bm{X}_{\rm spa}$, and at the forecasting horizon $\bm{X}_{t:t+H}$, respectively.
Figure 7: The dynamic (a) and static (b) adjacency matrices of weather dataset. The larger value indicates a closer relationship between nodes.
Figure 8: Visualization of relation strengths for three nodes located in different areas of Weather dataset. The orange triangles represent the selected target nodes. The darker color indicates a closer relationship between nodes. Panels (a), (b), and (c) show the dynamic relations of the nodes, while panels (d), (e), and (f) display the static relations.
Furthermore, we replace the temporal normalization modules to verify the effectiveness of SFL across various temporal normalization operations. We assessed SFL’s ability to preserve spatial distribution when combined with different temporal normalization modules: DAIN, Dish-TS, the Non-stationary Transformer and ST-norm. DAIN utilizes MLPs and a gate mechanism to learn the adaptive mean and standard deviation of input time series. The Non-stationary Transformer normalizes the time series and employs scaling factors to prevent removing excessive temporal information within the attention module. Dish-TS normalizes and de-normalizes the lookback horizon windows using different learned means and standard deviations. We applied temporal normalization after the frequency cutting operation, with SFL using the mean and standard deviation of the lookback windows to generate spatial factors. Given that RevIN performs normalization on feature dimensions, we conducted experiments applying only feature normalization to investigate whether this coarse-grained approach, which simultaneously normalizes both spatial and temporal distributions, is sufficient for spatio-temporal forecasting tasks. ST-norm applies spatial and temporal normalization to the inputs, enabling the model to better capture high-frequency spatial features and local temporal features. We compare three scenarios: applying T-norm only, applying ST-norm, and combining T-norm with the SFL module.
As shown in Table VII, SFL improves the performance of various temporal normalization methods, demonstrating its effectiveness as a general module for spatial distribution preservation. The models incorporating temporal operations and SFL outperform the model with RevIN. This finding suggests that normalization on feature dimensions alone is too coarse-grained and may not be adequate for distribution adaptation in spatio-temporal tasks. In ST-norm, the combination of spatial and temporal normalization improves performance. However, combining T-norm with the SFL module results in greater accuracy improvement, suggesting that the rescaling spatial distributions contribute more significantly to imitating propagation dynamics than normalization alone. Therefore, SFL after temporal normalization is an effective approach for distribution adaptation compared with instance-level and spatial and temporal level normalization.
V-C Dynamic and Static Relations Learning
In Fig. 7, we explore the adaptive process of our model by depicting the dynamic and static adjacency matrices within the DSFL module. Specifically, we showcase these matrices for the first input time step as an example. The dynamic adjacency matrix $\bm{A}_{\rm Dy}$ is derived from the similarity of time series between nodes according to Equation 6, while the static adjacency matrix is obtained by learning the static relations $\bm{A}_{\rm St}$ according to Equation 8. Fig. 7 illustrates the spatial relations between nodes of Weather dataset, with darker colors indicating stronger relationships. The dynamic adjacency matrix highlights the strength of relationships between nodes with similar signal patterns, whereas the static adjacency matrix focuses on the signals of individual nodes and some distributed nodes within the network. The contrast in focus enables the model to learn complementary features from both dynamic and static perspectives.
To clarify the differences between the learned dynamic and static relations for specific nodes, we select node $i$ from various locations and visualize the relationships between the selected node and other nodes, represented as $\bm{A}_{\rm{St},\textit{i}}$ and $\bm{A}_{\rm{Dy},\textit{i}}$. As shown in Fig. 8, subfigures (a), (b), and (c) depict the dynamic relations between target nodes and other nodes in the Weather dataset, while subfigures (d), (e), and (f) illustrate the static relations. The dynamic relations learned by DSFL concentrate around the target nodes, highlighting the significance of local connections. In contrast, the static relations reflect interactions between target nodes and distant nodes, indicating that the static adjacency matrix captures non-local relationships. This demonstrates that DSFL effectively learns comprehensive and non-overlapping spatial relations.
Moreover, the gate mechanism enables adaptive integration of dynamic and static features. We visualize the first channel of the gate signal $z$ from three layers of DSFL modules in Fig. 9, illustrating fusion ratio variations on the NYCBike2 test set. Each gate mechanism performs hierarchical fusion, where the blue-filled area represents the contribution of static features to the final deterministic features $\bm{X_{\rm D}}$, and the orange area represents the dynamic features. The fusion ratios vary in sync with the fluctuations of the input inflow and outflow series, demonstrating the gate mechanism’s ability to adaptively balance dynamic and static information.
Figure 9: Gate fusion ratio variations of DSFL modules on the NYCBike2 dataset during the test period. Panels (a), (b), and (c) show the first-channel dynamic and static fusion ratios for three DSFL modules, while panel (d) presents the corresponding input time series of the consistent gate ratios.
V-D Ablation Studies
To evaluate the effectiveness of each module in our network, we conduct an ablation study by systematically removing specific components. Specifically, we remove the DSFL module, the gate mechanism in DSFL module, the SFL module, the entire distribution adaptive module, and the Stochastic Learner to observe the resulting changes in prediction accuracy. The ablation strategies are detailed as follows:
  • w/o Sto: We remove the Stochastic Learner and used only the deterministic features for prediction. In this case, only $\bm{X}_{\rm D}$ is input into the decoder.
  • w/o Sta: We remove all modules related to non-stationarity and distribution adaptation, including normalization and de-normalization operations and the SFL module, and replace the de-stationary attention with standard attention [79].
  • w/o SFL: We remove the SFL module, resulting in a temporal-only normalization similar to that in the non-stationary Transformer.
  • w/o DSFL: We remove the DSFL module and use spatial attention instead.
  • w/o Gate: We remove the gate mechanism and replace it with a Linear layer mapping concatenated feature ${\rm F}_{\rm Cat}\in\mathbb{R}^{L\times N\times 2C}$ to shape $\mathbb{R}^{L\times N\times C}$.
The results of the ablation study are depicted in Fig. 10. The final model, incorporating all modules, achieves the best performance. It is evident that the SFL module contributes most significantly to improving prediction accuracy across different tasks. Removing all distribution adaptation modules (w/o Sta) has less impact compared to removing the SFL module (w/o SFL), highlighting SFL as a crucial and indispensable component for spatio-temporal distribution adaptation. The effectiveness of the Stochastic Learner varies depending on the task, aligning with the fact that task uncertainties differ. Comparing the experimental results of "w/o Gate" and "w/o DSFL," the prediction error increases more significantly when the gate mechanism in the DSFL module is removed than when the entire DSFL module is removed. This suggests that fusing dynamic and static representations with a fixed ratio, without considering the time-varying relationships between them, hinders accurate prediction and results in an unsuitable feature combination.
Figure 10: The results of ablation studies. Panels (a)–(c) show the MAE for inflow prediction, and panels (d)–(f) for outflow prediction, across the NYCBike1, NYCBike2, and NYCTaxi datasets, respectively. Each bar color denotes the removal of a specific module, as indicated in the legend. “Ours” represents the full model.
Moreover, our model outperforms the scenario where all normalization operations are removed, highlighting the necessity of normalization in spatio-temporal data processing.
VI Discussion and Conclusions
From the carried out experiments we can conclude the following:
  • The proposed method performs better in the spatio-temporal forecasting task compared with baseline methods at the cost of a moderate computation.
  • The proposed SFL module can be inserted with other temporal normalization methods and architectures, adapting distribution shifts in spatio-temporal context.
  • The proposed DSFL module is effective to both capture the dynamic and static spatial relations.
  • Each component (SFL, DSFL and Stochatic learner) in our DRAN model is provides value in prediction accuracy improvements. The adaptive fusion ratio derived from the gate mechanism is important for the integration of static and dynamic features.
Despite its strengths in accuracy performance, the methods suffers in terms of memory requirements and computational resources needs from scalability in large datasets or applications characterized by real-world systems. Additionally, the current design focuses on regular spatio-temporal patterns, making it less effective in scenarios characterized by the presence of abrupt changes or rare events.
To overcome above limitations, future work should focus on:
  • Scalability: Given the time costs associated with model training and inference, further research should investigate a more lightweight framework to deal with distribution shifts. This approach would improve scalability and enable the model to be effectively applied to large real-world datasets.
  • Adaptability and Transferability: While our framework currently emphasizes relation and distribution adaptation, it lacks mechanisms for dynamically adjusting network parameters based on learned knowledge and new inputs. Future work will focus on developing strategies to learn and update network parameters, drawing inspiration from techniques like EAST-Net [67], which generates sequence-specific, on-the-fly parameters. Enhancing adaptability will enable the model to better detect and respond to sudden changes and events.
VII Conclusion
To conclude, spatio-temporal forecasting is essential for understanding the states of complex systems, yet accurate predictions are often hindered by the dynamic and intricate nature of these systems. This study addresses the challenge of adapting to dynamic changes in spatio-temporal systems using neural networks. A DRAN framework is proposed to accommodate changes in distribution shifts, relations, and stochastic variations. Our approach includes a SFL to enable effective temporal normalization for spatio-temporal contexts. Additionally, we develop a DSFL to capture features from both dynamic and static relations. Furthermore, our framework enables to learn the deterministic and stochastic representations of features. Experimental results demonstrate the superiority of our method and the effectiveness of its components.

### v3

@@
 Normalization and de-normalization mechanisms are commonly employed to address distribution shifts in time series forecasting, particularly in scenarios that show temporal variation without strong spatial interactions [22, 23, 24, 25].
-Normalization is applied to the input data to align its distribution with that of the training set, while de-normalization maps the processed results back to the original temporal distribution.
-However, above methods, which apply normalization and de-normalization on instance [22] or temporal dimensions [25], may not be well-suited for processing spatio-temporal data using graph neural networks that rely on relational inductive bias.
+Normalization is applied to the input data to align its distribution with that of the training set, while de-normalization maps the processed results back to the original temporal distribution.However, above methods, which apply normalization and de-normalization on instance [22] or temporal dimensions [25], may not be well-suited for processing spatio-temporal data using graph neural networks that rely on relational inductive bias.
 In fact, temporal normalization scales the time series of nodes using different scaling factors as the temporal content evolves with time variance, hence causing erroneous scaled spatial messages propagating between nodes.
@@
 However, these approaches often overlook the dynamic interplay between static and dynamic spatial components, relying on fixed fusion schemes.
-This can lead to suboptimal spatio-temporal representations, especially in scenarios where the relative importance of static and dynamic features varies over time.
-Therefore, there is a pressing need for methods capable of adapting to both static and dynamic relations.
+This can lead to suboptimal spatio-temporal representations, especially in scenarios where the relative importance of static and dynamic features varies over time.Therefore, there is a pressing need for methods capable of adapting to both static and dynamic relations.
 In addition, spatio-temporal systems are inherently affected by various sources of uncertainty, including data inaccuracies, sensor noise, and sudden stochastic events that might change with time.
@@
 4:        Compute $\mu_{\bm{X}}$, $\sigma_{\bm{X}}$
-5:        Obtain $\bm{\textit{X}}_{\rm Norm}$ via Eq (1)
-6:        $\bm{\textit{X}}_{\rm tem}\xleftarrow{}\rm{DestationaryAtt}(\bm{\textit{X}}_{\rm Norm},\mu_{\bm{X}},\sigma_{\bm{X}})$ via Eq (2-4)
+5:        Obtain $\bm{\textit{X}}_{\rm Norm}$ via Eq. (1)
+6:        $\bm{\textit{X}}_{\rm tem}\xleftarrow{}\rm{DestationaryAtt}(\bm{\textit{X}}_{\rm Norm},\mu_{\bm{X}},\sigma_{\bm{X}})$ via Eq. (2-4)
 7:        $\mu_{\rm spa}$, $\sigma_{\rm spa}\xleftarrow{}\rm{SFL}(\bm{\textit{X}},\bm{\textit{X}}_{\rm tem},\mu_{\bm{X}},\sigma_{\bm{X}})$
-8:        Rescale according to Eq (5) and obtain $\bm{\textit{X}}_{\rm spa}$
-9:        $\rm{\textit{X}}_{\rm D}\xleftarrow{}\rm{DSFL}(\bm{\textit{X}}_{\rm spa},\bm{\textit{E}}_{\rm a})$ via Eq (6-11)
-10:        $\rm{\textit{X}}_{\rm S},\rm{\textit{X}}_{\rm rec}\xleftarrow{}\rm{StochasticLearner}(\bm{\textit{X}}_{\rm D})$ via Eq (12-15)
-11:        Compute $\mathcal{L}$ via Eq (17) and optimize $\theta$
+8:        Rescale according to Eq. (5) and obtain $\bm{\textit{X}}_{\rm spa}$
+9:        $\rm{\textit{X}}_{\rm D}\xleftarrow{}\rm{DSFL}(\bm{\textit{X}}_{\rm spa},\bm{\textit{E}}_{\rm a})$ via Eq. (6-11)
+10:        $\rm{\textit{X}}_{\rm S},\rm{\textit{X}}_{\rm rec}\xleftarrow{}\rm{StochasticLearner}(\bm{\textit{X}}_{\rm D})$ via Eq. (12-15)
+11:        Compute $\mathcal{L}$ via Eq. (17) and optimize $\theta$
 12:     end for
@@
 We generate the gate signal $\bm{z}$ by applying a FC layer to map $\bm{X}_{\rm Cat}$ from $\mathbbm{R}^{L\times N\times 2C}$ to $\mathbbm{R}^{L\times N\times C}$, followed by a Sigmoid activation function; here, FC denotes the fully connected layer.
-Finally, we employ Equation (12) to fuse the dynamic and static features, producing the final deterministic features $\bm{X}_{\rm D}$, which represent the final deterministic features learned from the input time series.
+Finally, we employ Eq.
+(12) to fuse the dynamic and static features, producing the final deterministic features $\bm{X}_{\rm D}$, which represent the final deterministic features learned from the input time series.
 TABLE I: Datasets details
@@
 It fine-tunes the static graph using node similarities, effectively fusing dynamic and static features.
-TESTAM employs a mixture-of-experts model with three experts: one for temporal modeling, one for spatio-temporal modeling with a static graph, and one for spatio-temporal dependency modeling with a dynamic graph.
-We also fine-tune the hyperparameters of baseline methods to ensure fairness, especially for datasets not included in their official implementations.
+TESTAM employs a mixture-of-experts model with three experts: one for temporal modeling, one for spatio-temporal modeling with a static graph, and one for spatio-temporal dependency modeling with a dynamic graph.We also fine-tune the hyperparameters of baseline methods to ensure fairness, especially for datasets not included in their official implementations.
 We evaluate the performance of these models using two metrics: MAE and Mean Absolute Percentage Error (MAPE).
@@
 Specifically, we showcase these matrices for the first input time step as an example.
-The dynamic adjacency matrix $\bm{A}_{\rm Dy}$ is derived from the similarity of time series between nodes according to Equation 6, while the static adjacency matrix is obtained by learning the static relations $\bm{A}_{\rm St}$ according to Equation 8.
+The dynamic adjacency matrix $\bm{A}_{\rm Dy}$ is derived from the similarity of time series between nodes according to Eq.
+6, while the static adjacency matrix is obtained by learning the static relations $\bm{A}_{\rm St}$ according to Eq.
+8.
 Fig.